A sample of size 3 is taken from a population of the number of students in each class of Grade 9 students listed as 25, 30, 35, and 40.



Compare the population mean and the sampling distribution of the sample means.




1
Expert's answer
2022-03-16T05:25:49-0400

Solution

Population mean= (25+30+35+40)/4

     =32.5

4C3=4

Sample means

(25,30,35)  

Mean=(25,30,35)/3=30

(25,30,40)

Mean =(25+30+40)/3=31.6667

(25,35,40)

Mean= (25+35+40)/3=33.333

(30,35,40)

Mean= (30+35+40)/3=35

(30+31.6667+33.333+35)/4= 32.5

The population mean and the sampling distribution of sample means are equal.